OpenStudy (anonymous):

jane works north of her home. her husband Bob works east. they leave for work at the same time. By the time Jane is 7 miles from home, the distance between them is one mile more than Bobs distance from home. How far from home is bob?

OpenStudy (anonymous):

theorem of pythagoras i would say

OpenStudy (anonymous):

7^2 + (x)^2 = (x+1)^2

OpenStudy (anonymous):

this produces the quadratic \[49 + x ^{2} = x^2 + 2x + 1\]

OpenStudy (anonymous):

x=24

OpenStudy (anonymous):

i dont understand quadratics =(

OpenStudy (anonymous):

u dont have to in this case as you may simply subtract x^2 from both sides and you get a linear equation in x which is then solved by subtracting one from both sides and dividing by 2

OpenStudy (anonymous):

i dont understand any of it =(

OpenStudy (anonymous):

pythagoras theorem states \[hypthenuse^2 = adjacent^2 + opposite^2\]

OpenStudy (anonymous):

can u see how their movements form a right angled triangle? draw it for urself

OpenStudy (anonymous):

i thought it was a +b=c

OpenStudy (anonymous):

pythagoras' theorem? no the sides are squared

OpenStudy (anonymous):

i dont understand. sorry

OpenStudy (anonymous):

start by drawing the problem. mark a dot on the page for their home and then take a line 7 north for janes movement.

OpenStudy (anonymous):

now go back to the dot and take a line east for the dot (length doesnt matter) let this line have length x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now mark the positions of mary and bob by a dot. now join those two dots and let the length of this new line be x+1 because the question says the distance between them is one more than how far bob is from home

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now can you see the right angled triangle? the hypothenuse (longest side) has length x+1. apply pythagoras theorem to get \[7^2 + (x)^2 = (x+1)^2\]

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

7x to the 4th = 2x+2?? i dont know what im doing...im far from help

OpenStudy (anonymous):

i think you are confused between powers of a number and mulitplication\[(x+1)^2 = (x+1)(x+1) = x^2 +x + x +1 = x^2 +2x +1\]

OpenStudy (anonymous):

which is the right hand side, left hand side is just 49 + x^2

OpenStudy (anonymous):

omg 49 +2x=2x??

OpenStudy (anonymous):

almost \[49+x^2 = x^2 + 2x + 1 \rightarrow 49 = 2x +1 \rightarrow 2x=48 \]

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

verify the steps yourself to make sure u understand

OpenStudy (anonymous):

i dont

OpenStudy (anonymous):

49+x times x=(x+1)(x+1)

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

only i dont know how to answer that 49x2=x2+1x+1?

OpenStudy (anonymous):

thank you for your help